# Dynamic Rank Reinforcement Learning for Adaptive Low-Rank Multi-Head Self Attention in Large Language Models

### v1

@@ Abstract
We propose Dynamic Rank Reinforcement Learning (DR-RL), a novel framework that adaptively optimizes the low-rank factorization of Multi-Head Self-Attention (MHSA) in Large Language Models (LLMs) through the integration of reinforcement learning and online matrix perturbation theory. While traditional low-rank approximations often rely on static rank assumptions—limiting their flexibility across diverse input contexts—our method dynamically selects ranks based on real-time sequence dynamics, layer-specific sensitivities, and hardware constraints. The core innovation lies in an RL agent that formulates rank selection as a sequential policy optimization problem, where the reward function strictly balances attention fidelity against computational latency. Crucially, we employ online matrix perturbation bounds to enable incremental rank updates, thereby avoiding the prohibitive cost of full decomposition during inference. Furthermore, the integration of a lightweight Transformer-based policy network and batched Singular Value Decomposition (SVD) operations ensures scalable deployment on modern GPU architectures. Experiments demonstrate that DR-RL maintains downstream accuracy statistically equivalent to full-rank attention while significantly reducing Floating Point Operations (FLOPs), particularly in long-sequence regimes ($L>4096$). This work bridges the gap between adaptive efficiency and theoretical rigor in MHSA, offering a principled, mathematically grounded alternative to heuristic rank reduction techniques in resource-constrained deep learning.  [^1] Source code and experiment logs are available at: https://github.com/canererden/DR_RL_Project.
I Introduction
Large Language Models (LLMs) have revolutionized natural language processing by capturing complex linguistic patterns and generating human-like text with unprecedented fidelity. At the heart of these models lies the Multi-Head Self-Attention (MHSA) mechanism, which enables the parallel processing of diverse contextual relationships within input sequences [vaswani2017attention]. While MHSA provides remarkable expressive power, its computational complexity scales quadratically with respect to sequence length ($\mathcal{O}(N^{2})$), constituting a significant bottleneck for scalability, particularly when processing long documents or deploying models on resource-constrained edge devices.
@@ I Introduction
To mitigate the computational burden of MHSA, several approximation techniques have been proposed, with low-rank factorization emerging as a promising direction. By approximating dense attention matrices using lower-rank representations, these methods reduce memory footprint and accelerate computation [sainath2013low]. However, existing techniques typically employ static rank selection strategies that remain fixed during inference. This static approach overlooks the dynamic nature of linguistic structures; the optimal rank required to capture semantic dependencies may vary substantially across different layers, input sequences, and even individual attention heads. Consequently, a “one-size-fits-all” rank assignment often leads to a suboptimal Pareto frontier: either incurring excessive computation for simple tokens or suffering information loss in complex contexts.
We address this limitation by introducing a Dynamic Rank Reinforcement Learning (DR-RL) framework that integrates deep reinforcement learning with online matrix perturbation theory. The key insight is that rank selection can be formulated as a sequential decision-making problem, where an RL agent learns a policy to adjust ranks based on the evolving characteristics of the input and the current latent state of the model [sutton2018reinforcement]. This approach differs fundamentally from prior work in three aspects:
@@
Unlike heuristic pruning or fixed compression methods, DR-RL allows the model to allocate computational resources adaptively, assigning higher ranks to semantically dense regions and lower ranks to redundant ones. The integration of perturbation-based updates specifically addresses the latency concerns of dynamic architectures, avoiding the prohibitive overhead of exhaustive rank search.
The main contributions of this paper are summarized as follows:
@@
  • Theoretical Grounding: We derive bounds based on Matrix Perturbation Theory [niklasson2004density] to guide the RL agent, ensuring that rank modifications remain within a stability region that preserves semantic integrity.
@@ II Related Work
This section situates our work at the intersection of low-rank approximations for attention mechanisms, reinforcement learning for neural architecture search, and matrix perturbation theory.
II-A Low-Rank Approximations for Attention Mechanisms
The computational bottleneck of standard Multi-Head Self-Attention (MHSA), characterized by $\mathcal{O}(N^{2})$ complexity, has necessitated the development of efficient approximation techniques. Linformer [wang2020linformer] introduced a projection-based approach, mapping key and value matrices to a fixed low-dimensional space to achieve linear complexity while retaining performance. Similarly, Performer [choromanski2020rethinking] utilized orthogonal random features to approximate the softmax kernel, enabling scalable attention without explicit matrix instantiation. While these methods demonstrate that low-rank structures can preserve semantic information, they fundamentally rely on static rank assumptions. The rank is determined prior to inference and remains constant across varying input densities and layer depths, potentially creating a rigidity that fails to capture the dynamic complexity of diverse linguistic contexts.
More recently, research has shifted towards flexible adaptation. DyLoRA [valipour2023dylora] proposed a dynamic low-rank adaptation strategy for parameter-efficient fine-tuning (PEFT), eliminating the need for exhaustive rank search during training. However, while DyLoRA addresses training efficiency, it does not solve the challenge of input-dependent rank selection during inference. Other works, such as the continual low-rank scaled dot-product attention framework [picon2024continual], focus on mitigating catastrophic forgetting in continual learning scenarios rather than optimizing inference latency. Unlike these approaches, our framework targets the real-time, context-aware modulation of rank during the forward pass, distinct from continual learning or static compression strategies.
II-B Reinforcement Learning for Neural Architecture Adaptation
@@
Reinforcement Learning (RL) has been successfully applied to optimize discrete architectural decisions. Prior works have utilized RL for hyperparameter tuning [zoph2016neural], layer selection [wu2018blockdrop], and dynamic computation allocation. These approaches share our perspective of formulating architectural configuration as a sequential decision-making problem. For instance, ALoRE [du2024alore] demonstrated the efficacy of aggregating low-rank experts for visual domain adaptation. Similarly, adaptive computation in Vision Transformers [dong2024lowrank] has been explored, though primarily through token pruning or scaling factors rather than spectral modulation. While methods like BlockDrop [wu2018blockdrop] demonstrate the efficacy of skipping residual blocks to save computation, they operate at a coarse granularity.
Crucially, none of these existing methods address the specific challenge of spectral rank selection in attention matrices. This problem is distinct because the decision space involves a trade-off between the algebraic fidelity of the matrix approximation and computational throughput (FLOPs), requiring a reward function that explicitly balances numerical precision with system latency.
II-C Matrix Perturbation Theory in Deep Learning
Matrix perturbation theory provides the analytical tools to quantify how variations in matrix components affect spectral properties. Recent studies have established perturbation bounds for various neural components to ensure robustness [zhang2019robust, niklasson2004density]. In the context of attention mechanisms, perturbation analysis has primarily been employed to study the stability of Transformer representations against input noise or adversarial attacks [zhang2020accelerating].
Our work extends this theoretical domain by applying perturbation bounds specifically to structural rank modifications ($\mathbf{A}\rightarrow\mathbf{A}_{r_{t}}$). Rather than analyzing external noise, we derive bounds for the internal approximation error induced by rank reduction. This allows us to theoretically guarantee that the RL agent’s actions remain within a “safe” trust region, bridging the gap between heuristic compression and theoretically grounded optimization.
To summarize, our proposed DR-RL framework synthesizes insights from these diverse strands. Unlike fixed-rank approximations [wang2020linformer, choromanski2020rethinking], we enable granularity through input-dependent selection. Compared to general Neural Architecture Search (NAS) [zoph2016neural], we incorporate domain-specific constraints from linear algebra. Finally, relative to existing adaptive methods [valipour2023dylora], [graves2016adaptive] we provide rigorous theoretical guarantees via perturbation theory, offering a significant advance in the design of robust and efficient attention mechanisms.
III Preliminaries: Low-Rank MHSA and Online Matrix Perturbation
@@
III-A Low-Rank Multi-Head Self-Attention
@@
where $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{n\times d}$ denote the query and key matrices, respectively. The computational bottleneck arises from the $\mathbf{Q}\mathbf{K}^{T}$ multiplication and the subsequent operations on the $n\times n$ matrix, leading to $\mathcal{O}(n^{2})$ complexity.
Low-rank approximations mitigate this by factorizing the attention matrix into products of lower-rank matrices. Let $\mathbf{A}\approx\mathbf{U}\mathbf{V}^{T}$, where $\mathbf{U},\mathbf{V}\in\mathbb{R}^{n\times r}$ with rank $r\ll n$. This formulation potentially reduces the computational complexity from $\mathcal{O}(n^{2}d)$ to $\mathcal{O}(nrd)$, preserving the model’s expressive power provided that the intrinsic rank $r$ is chosen appropriately [wang2020linformer]. The fidelity of this approximation depends critically on the selection of $r$, which should ideally adapt to the input characteristics ($\mathbf{x}$) and layer depth ($l$).
III-B Singular Value Decomposition for Attention Approximation
@@
III-C Online Matrix Perturbation Theory
@@ III-D Computational Considerations
IV Methodology: Dynamic Rank Selection via Reinforcement Learning and Perturbation Analysis
@@ IV-A Problem Formulation as MDP
IV-A1 State Space ($\mathcal{S}$)
@@
IV-A2 Action Space ($\mathcal{A}$)
@@
IV-A3 Policy Network ($\pi_{\theta}$)
@@
IV-A4 Reward Function ($\mathcal{R}$)
@@
IV-B Online Matrix Perturbation for Real-Time Adaptation
@@
IV-C Integration of RL and Perturbation Theory
@@ IV-C2 Incremental SVD Update
This reduces computational overhead by avoiding full re-decomposition, providing a speedup proportional to $(r^{\prime}-r)/r^{\prime}$.
Finally, we augment the reward function with a stability penalty derived from the perturbation norm:
@@
IV-D Context-Aware Projection and Spectral Energy
@@
IV-E Implementation Strategy
IV-E1 Policy Network Architecture
@@
IV-E2 Batched Execution and Power Iteration
@@
IV-E3 Hybrid Training
We employ a warm-start strategy where the policy is first pretrained via Behavior Cloning (BC) on trajectories generated by an offline Oracle (greedy search). Subsequently, the agent is fine-tuned using the PPO algorithm [sutton2018reinforcement] with the reward function defined in Eq. 13.
V Experimental Evaluation
V-A Experimental Setup
To evaluate the efficacy of the proposed Dynamic Rank RL (DR-RL) framework, we conducted extensive experiments on standard language modeling benchmarks. We implemented our approach on top of the standard Transformer decoder architecture [vaswani2017attention] and compared it against the following baselines:
  • Full-Rank Attention: Standard Multi-Head Self-Attention (MHSA) without any low-rank approximation, serving as the upper bound for performance.
  • Fixed Low-Rank: MHSA with a static rank selection ($r=32$) fixed across all layers and sequences, representative of static compression methods [wang2020linformer].
  • Adaptive SVD: A heuristic approach that dynamically selects ranks based on a cumulative singular value energy threshold (e.g., maintaining 90% variance) [yang2021adaptive].

### v2

# Dynamic Rank Reinforcement Learning for Adaptive Low-Rank Multi-Head Self-Attention in Large Language Models

@@ Abstract
Dynamic Rank Reinforcement Learning (DR-RL) approximations rely on static rank assumptions, limiting their flexibility across diverse linguistic contexts. Our method dynamically modulates ranks based on real-time sequence dynamics, layer-specific sensitivities, and hardware constraints. The core innovation is a deep reinforcement learning agent that formulates rank selection as a sequential policy optimization problem, strictly balancing attention fidelity against computational latency. To ensure stability during inference, we derive and employ online matrix perturbation bounds, enabling incremental rank updates without the prohibitive cost of full decomposition. Furthermore, the integration of a lightweight Transformer-based policy network and batched Singular Value Decomposition (SVD) operations ensures scalable deployment on modern architectures. Extensive experiments demonstrate that DR-RL significantly reduces Floating Point Operations (FLOPs) by over 40% in long-sequence regimes ($L>4096$) while maintaining downstream accuracy statistically equivalent to full-rank attention. Beyond standard language modeling benchmarks, we validate the real-world applicability of DR-RL on the GLUE benchmark. Specifically, our method achieves 92.78% accuracy on the SST-2 sentiment analysis task, matching the performance of full-rank baselines and outperforming static low-rank methods, such as Performer and Nyströmformer, by a significant margin.
00footnotetext: Article information: International Journal of Complexity in Applied Science and Technology, 2026.
DOI: https://doi.org/10.1504/IJCAST.2026.309733
00footnotetext: Biographical statement: Dr. Caner Erden is an Associate Professor of Computer Engineering at Sakarya University of Applied Sciences (SUBÜ), Turkey, and a Senior Researcher at the AI Research and Application Center. He holds a Ph.D. from Sakarya University (2019). His research interests include data science, machine learning, operations research, and meta-heuristic algorithms.
Keywords: Large Language Models, LLM, Multi-Head Self-Attention, Reinforcement Learning, Low-Rank Approximation, Dynamic Rank Selection
1 Introduction
@@
Large Language Models (LLMs) have revolutionized natural language processing by capturing complex linguistic patterns and generating human-like text with unprecedented fidelity. At the heart of these models lies the Multi-Head Self-Attention (MHSA) mechanism, which enables the parallel processing of diverse contextual relationships within input sequences [28]. While MHSA provides remarkable expressive power, its computational complexity scales quadratically with respect to sequence length ($\mathcal{O}(N^{2})$), constituting a significant bottleneck for scalability, particularly when processing long documents or deploying models on resource-constrained edge devices. To mitigate the computational burden of MHSA, several approximation techniques have been proposed, with low-rank factorization emerging as a promising direction [8]. By approximating dense attention matrices using lower-rank representations, these methods reduce memory footprint and accelerate computation [21, 7, 12, 24]. However, existing techniques typically employ static rank selection strategies that remain fixed during inference. This static approach overlooks the dynamic nature of linguistic structures; the optimal rank required to capture semantic dependencies may vary substantially across different layers, input sequences, and even individual attention heads. Consequently, a “one-size-fits-all” rank assignment often leads to a suboptimal Pareto frontier: either incurring excessive computation for simple tokens or suffering information loss in complex contexts. We address this limitation by introducing a Dynamic Rank Reinforcement Learning (DR-RL) framework that integrates deep reinforcement learning with online matrix perturbation theory. The key insight is that rank selection can be formulated as a sequential decision-making problem, where an RL agent learns a policy to adjust ranks based on the evolving characteristics of the input and the current latent state of the model [26]. This approach differs fundamentally from prior work in three aspects:
@@
Unlike heuristic pruning or fixed compression methods, DR-RL allows the model to allocate computational resources adaptively, assigning higher ranks to semantically dense regions and lower ranks to redundant ones. The integration of perturbation-based updates specifically addresses the latency concerns of dynamic architectures, avoiding the prohibitive overhead of exhaustive rank search. The main contributions of this paper are summarized as follows:
@@
  • Theoretical Grounding: We derive bounds based on Matrix Perturbation Theory [18] to guide the RL agent, ensuring that rank modifications remain within a stability region that preserves semantic integrity.
@@ 2 Related Work
This section situates our work at the intersection of low-rank approximations for attention mechanisms, dynamic pruning strategies, reinforcement learning for neural architecture search, and matrix perturbation theory.
2.1 Low-Rank Approximations for Attention Mechanisms
The computational bottleneck of standard MMHSA, characterized by $\mathcal{O}(N^{2})$ complexity, has necessitated the development of efficient approximation techniques. Linformer [29] introduced a projection-based approach, mapping key and value matrices to a fixed low-dimensional space to achieve linear complexity while retaining performance. Similarly, Performer [4] utilized orthogonal random features to approximate the softmax kernel, enabling scalable attention without explicit matrix instantiation. While these methods demonstrate that low-rank structures can preserve semantic information, they fundamentally rely on static rank assumptions. The rank is determined prior to inference and remains constant across varying input densities and layer depths, potentially creating a rigidity that fails to capture the dynamic complexity of diverse linguistic contexts. Alternative spectral approaches, such as DCT-Former [22], attempt to reduce memory usage by approximating self-attention via Discrete Cosine Transforms. However, this necessitates modifying the fundamental kernel operation, whereas DR-RL maintains the standard Multi-Head Attention mechanism, learning per-head ranks to ensure broader compatibility with pre-trained models. Recent approaches have sought to mitigate this rigidity through more sophisticated architectural adjustments. For instance, Qin et al. [20] introduced Dynamic Bilinear Low-Rank Attention (DBA), which achieves linear complexity via dynamic projections; however, unlike our DR-RL, it relies on a fixed low-rank structure per layer rather than an adaptive one. Similarly, Vicinity Attention [25] incorporates 2D locality for efficiency in vision transformers but fixes the feature reduction blocks architecturally. Comprehensive surveys on LLM compression and adaptation underscore the dominance of static methods in the current landscape. As categorized by Mao et al. [13], most Low-Rank Adaptation (LoRA) variants focus on optimizing fixed-rank updates for parameter-efficient fine-tuning rather than inference-time flexibility. Similarly, systematic reviews by Kim et al. [10] and Zhu et al. [36] highlight that while quantization and pruning are widely adopted for model compression, they typically result in “one-shot” compressed models that lack the ability to adaptively modulate their computational budget per input. DR-RL fills this specific gap identified in the literature by offering an online, RL-driven rank modulation mechanism within standard Transformer layers. More recently, research has shifted towards flexible adaptation. DyLoRA [27] proposed a dynamic low-rank adaptation strategy for parameter-efficient fine-tuning (PEFT), eliminating the need for exhaustive rank search during training. However, while DyLoRA addresses training efficiency, it does not solve the challenge of input-dependent rank selection during inference. Other works, such as the continual low-rank scaled dot-product attention framework [19], focus on mitigating catastrophic forgetting in continual learning scenarios rather than optimizing inference latency. Unlike these approaches, our framework targets the real-time, context-aware modulation of rank during the forward pass.
2.2 Dynamic Pruning and Sparse Attention
Another prevalent direction for reducing computational overhead is dynamic pruning, which selectively discards redundant information. Several recent studies have proposed adaptive mechanisms to reduce the token or head count during inference. For example, Zheng et al. [35] combined linearized attention with a parameter-free adaptive token pruning scheme to reduce FLOPs. Similarly, Ahmadpanah et al. [1] leveraged task-specific attention patterns to prune tokens dynamically in LLMs based on adaptive thresholds. In the domain of computer vision, Nawaz et al. [17] applied dynamic token pruning for facial emotion recognition, while Messaoud et al. [15] proposed adaptive head pruning to eliminate less informative attention heads entirely. While effective, these methods typically involve binary decisions (keep/drop) that may result in irreversible information loss. In contrast, DR-RL preserves all tokens and heads but dynamically modulates the spectral rank of the attention matrix via reinforcement learning, offering a finer-grained, continuous control over the fidelity-efficiency trade-off.
2.3 Reinforcement Learning for Neural Architecture Adaptation
@@
RL has been successfully applied to optimize discrete architectural decisions. Prior works have utilized RL for hyperparameter tuning [38], layer selection [30], and dynamic computation allocation. These approaches share our perspective of formulating architectural configuration as a sequential decision-making problem. For instance, ALoRE [6] demonstrated the efficacy of aggregating low-rank experts for visual domain adaptation. Similarly, adaptive computation in Vision Transformers [5] has been explored, though primarily through token pruning or scaling factors rather than spectral modulation. While methods like BlockDrop [30] demonstrate the efficacy of skipping residual blocks to save computation, they operate at a coarse granularity. However, a fundamental distinction lies in the timing of the optimization. Most recent RL-based Neural Architecture Search (NAS) frameworks operate offline to discover a fixed optimal structure. For instance, Cassimon et al. [2] utilized scalable RL agents to iteratively search for optimal Transformer backbones, while Shi et al. [23] and Li et al. [11] applied similar offline RL strategies to optimize GANs and CNNs for specific domains like EEG analysis. In contrast to these design-time optimizations, DR-RL employs an RL agent online during inference. Instead of searching for a single static architecture, our agent continuously modulates the spectral rank of attention matrices based on the instantaneous complexity of the input sequence. Crucially, none of these existing methods address the specific challenge of spectral rank selection in attention matrices. This problem is distinct because the decision space involves a trade-off between the algebraic fidelity of the matrix approximation and computational throughput (FLOPs), requiring a reward function that explicitly balances numerical precision with system latency.
2.4 Matrix Perturbation Theory in Deep Learning
Matrix perturbation theory provides the analytical tools to quantify how variations in matrix components affect spectral properties. Recent studies have established perturbation bounds for various neural components to ensure robustness [33, 18]. In the context of attention mechanisms, perturbation analysis has primarily been employed to study the stability of Transformer representations against input noise or adversarial attacks [34]. Finally, broader surveys on transformer inference optimization [3] emphasize the growing need for hardware-aware efficiency. Recent hardware-centric studies, such as the work by Hsiung and Chang [9], have demonstrated the benefits of co-designing accelerators with dynamic pruning strategies. Our work complements this landscape by introducing dynamic rank adaptation as an orthogonal axis of optimization. Unlike hardware-specific pruning, DR-RL offers a mathematically grounded, algorithmic reduction in FLOPs that can be leveraged across diverse hardware backends, from commodity CPUs to enterprise GPUs. Our work extends this theoretical domain by applying perturbation bounds specifically to structural rank modifications ($\mathbf{A}\rightarrow\mathbf{A}_{r_{t}}$). Rather than analyzing external noise, we derive bounds for the internal approximation error induced by rank reduction. This allows us to theoretically guarantee that the RL agent’s actions remain within a “safe” trust region, bridging the gap between heuristic compression and theoretically grounded optimization. To summarize, our proposed DR-RL framework synthesizes insights from these diverse strands. Unlike fixed-rank approximations [29, 4], we enable granularity through input-dependent selection. Compared to general Neural Architecture Search (NAS) [38], we incorporate domain-specific constraints from linear algebra. Finally, relative to existing adaptive methods [27], we provide rigorous theoretical guarantees via perturbation theory, offering a significant advance in the design of robust and efficient attention mechanisms.
3 Preliminaries: Low-Rank MHSA and Online Matrix Perturbation
@@
3.1 Low-Rank Multi-Head Self-Attention
@@
where $\mathbf{Q},\mathbf{K}\in\mathbb{R}^{n\times d}$ denote the query and key matrices, respectively. The computational bottleneck arises from the $\mathbf{Q}\mathbf{K}^{T}$ multiplication and the subsequent operations on the $n\times n$ matrix, leading to $\mathcal{O}(n^{2})$ complexity. Low-rank approximations mitigate this by factorizing the attention matrix into products of lower-rank matrices. Let $\mathbf{A}\approx\mathbf{U}\mathbf{V}^{T}$, where $\mathbf{U},\mathbf{V}\in\mathbb{R}^{n\times r}$ with rank $r\ll n$. This formulation potentially reduces the computational complexity from $\mathcal{O}(n^{2}d)$ to $\mathcal{O}(nrd)$, preserving the model’s expressive power provided that the intrinsic rank $r$ is chosen appropriately [29]. The fidelity of this approximation depends critically on the selection of $r$, which should ideally adapt to the input characteristics ($\mathbf{x}$) and layer depth ($l$).
3.2 Singular Value Decomposition for Attention Approximation
@@
3.3 Online Matrix Perturbation Theory
@@ 3.4 Computational Considerations
4 Methodology: Dynamic Rank Selection via Reinforcement Learning and Perturbation Analysis
@@ 4.1 Problem Formulation as MDP
4.1.1 State Space ($\mathcal{S}$)
@@
4.1.2 Action Space ($\mathcal{A}$)
@@
4.1.3 Policy Network ($\pi_{\theta}$)
@@
4.1.4 Reward Function ($\mathcal{R}$)
@@
4.2 Online Matrix Perturbation for Real-Time Adaptation
@@
4.3 Integration of RL and Perturbation Theory
@@ 4.3.2 Incremental SVD Update
This reduces computational overhead by avoiding full re-decomposition, providing a speedup proportional to $(r^{\prime}-r)/r^{\prime}$. Finally, we augment the reward function with a stability penalty derived from the perturbation norm:
@@
4.4 Context-Aware Projection and Spectral Energy
@@
4.5 Implementation Strategy
4.5.1 Policy Network Architecture
@@
4.5.2 Batched Execution and Power Iteration
@@
4.5.3 Hybrid Training
We employ a warm-start strategy where the policy is first pretrained via Behavior Cloning (BC) on trajectories generated by an offline Oracle (greedy search). Subsequently, the agent is fine-tuned using the PPO algorithm [26] with the reward function defined in Eq. 13.
5 Experimental Evaluation
5.1 Experimental Setup
To evaluate the efficacy of the proposed Dynamic Rank RL (DR-RL) framework, we conducted extensive experiments on standard language modeling benchmarks. To demonstrate the computational efficiency and accessibility of our proposed DR-RL framework, all experiments were conducted on a commodity workstation rather than large-scale server clusters. The system was equipped with an Apple Silicon processor featuring 12 physical cores and 16 GB of unified memory. For hardware acceleration, we utilized the Metal Performance Shaders (MPS) backend on macOS (Darwin Kernel 25.2.0). The software environment relied on Python 3.12.12 and PyTorch 2.9.1. Optimization was performed using the AdamW optimizer with a linear learning rate scheduler. The successful training of the model on this resource-constrained setup empirically validates the low operational footprint of DR-RL compared to full-rank attention mechanisms that typically mandate high-end enterprise GPUs. We implemented our approach on top of the standard Transformer decoder architecture [28] and compared it against the following baselines:
  • Full-Rank Attention: Standard MHSA without any low-rank approximation, serving as the upper bound for performance.
  • Fixed Low-Rank: MHSA with a static rank selection ($r=32$) fixed across all layers and sequences, representative of static compression methods [29].
  • Adaptive SVD: A heuristic approach that dynamically selects ranks based on a cumulative singular value energy threshold (e.g., maintaining 90% variance) [32].